OpenStudy (anonymous):

An aqueous solution freezes at -3.47⁰C. What is its boiling point? Show all calculations leading to an answer.

OpenStudy (aaronq):

what are the formulas for freezing point depression and boiling point elevation?

OpenStudy (anonymous):

freezing point depression is ΔT = iK f m boiling point elevation is ΔTb = Kbm

OpenStudy (aaronq):

Okay, so use the first one to find molality, m then use that value in the second equation

OpenStudy (anonymous):

how do I find molality? sorry I'm very confused

OpenStudy (aaronq):

In the first formula you posted (freezing point depression), plug the values in for \(\Delta T\) is given in the question, you know \(K_f\) because it's a constant. assume that i is 1

OpenStudy (anonymous):

ok so -3.47 C = (1)(1.86)(m) is this correct so far?

OpenStudy (aaronq):

yep that's right

OpenStudy (anonymous):

ok so -3.47 / 1.86 = -2.38 yes? whats after that

OpenStudy (aaronq):

-1.86 should be negative, so you end up with a positive value m=2.38 now use this value in the second equation (with the constant Kb)

OpenStudy (anonymous):

ok so 0.512 x 2.38 = 1.22 is this right?

OpenStudy (aaronq):

yep, that's the change in temperature from the normal boiling point of water that's not the final answer though, re-read the question

OpenStudy (anonymous):

so how do I find the boiling point mentioned in the question? do I add it to the freezing point?

OpenStudy (aaronq):

nope, you add it to the normal boiling point of water

OpenStudy (anonymous):

so 1.22 + 100 = 101.22?

OpenStudy (aaronq):

yep

OpenStudy (anonymous):

Thats it? Thank you so much for your help :)

OpenStudy (aaronq):

yes, thats the answer. no problem !